# WeatherGFM: Learning A Weather Generalist Foundation Model via In-context Learning

### v1

@@ Abstract
The Earth’s weather system encompasses intricate weather data modalities and diverse weather understanding tasks, which hold significant value to human life. Existing data-driven models focus on single weather understanding tasks (e.g., weather forecasting). Although these models have achieved promising results, they fail to tackle various complex tasks within a single and unified model. Moreover, the paradigm that relies on limited real observations for a single scenario hinders the model’s performance upper bound. In response to these limitations, we draw inspiration from the in-context learning paradigm employed in state-of-the-art visual foundation models and large language models. In this paper, we introduce the first generalist weather foundation model (WeatherGFM), designed to address a wide spectrum of weather understanding tasks in a unified manner. More specifically, we initially unify the representation and definition of the diverse weather understanding tasks. Subsequently, we devised weather prompt formats to manage different weather data modalities, namely single, multiple, and temporal modalities. Finally, we adopt a visual prompting question-answering paradigm for the training of unified weather understanding tasks. Extensive experiments indicate that our WeatherGFM can effectively handle up to ten weather understanding tasks, including weather forecasting, super-resolution, weather image translation, and post-processing. Our method also showcases generalization ability on unseen tasks.
@@ 1 Introduction
Modeling Earth weather systems involves a series of complex subprocesses that are intended to transform intricate Earth observation data into applications like weather forecasting (Chen et al., 2023a; Bi et al., 2023a), downscaling (Chen et al., 2022; Xu et al., 2024; Ling et al., 2024; Liu et al., 2024c), assimilation (Huang et al., 2024; Wang et al., 2024b), retrieval (Liu et al., 2011), and bias correction (Gong et al., 2024a; b). During the past decade, many data-driven methods have been investigated for various weather understanding tasks and delivering desirable performance on specific tasks. For example, recent studies using large-scale data (e.g., ERA5 reanalysis data (Hersbach et al., 2020)) have exceeded the accuracy of conventional numerical weather forecasts. However, current weather foundational models face challenges regarding generalizability and data scale limitations. On the one hand, the Earth observation system consists of a variety of observation devices, such as satellites, radar, and weather stations, which produce diverse modalities of data. Consequently, designing a specific model for a single-task scenario is highly complex, time-consuming, and labor-intensive. On the other hand, large-scale data in fields such as computer vision can be obtained at a low cost, whereas weather understanding tasks face an intrinsic bottleneck in data scale due to restrictions on individual scenes and single observation devices as shown in Table 1. For instance, local short-term precipitation forecasting models can only utilize a finite range of observational data.
A significant trend in AI research is the development of foundation models, shifting towards large-scale pre-training and in-context learning. This paradigm enables unified processing of a multitude of complex tasks and generalization to unseen tasks. For example, large language models (LLMs) can perform a variety of language-centric tasks (e.g., question answering, machine translation and dialogue state tracking (Feng et al., 2023)) by combining language input-output examples with new query inputs (prompts) without optimizing model parameters (Brown, 2020). Similarly, vision foundation models (Wang et al., 2023b; Chen et al., 2024b) employ visual prompts with query inputs to carry out diverse image-centric tasks, such as semantic segmentation, depth estimation, and image restoration. These studies highlight the significant potential of generalist foundational models.
@@
Computer Vision HAT (Chen et al., 2023d) Low-cost Image super-resolution (SR)
@@
Cascast (Gong et al., 2024a) High-cost Post-processing
Climax (Nguyen et al., 2023) High-cost Weather forecasting, Weather image SR Requires fine-tuning
@@
In this paper, we first propose a weather generalist foundation model, WeatherGFM, to uniformly address a variety of complex weather understanding tasks and data modalities. Unlike prior studies that focused on weather forecasting, our proposed method can expand the task scope to weather forecasting, weather super-resolution (i.e., weather downscaling), weather image translation (i.e., synthetic weather radar) (Veillette et al., 2020), and post-processing (Gong et al., 2024a). These tasks all belong to the domain of weather understanding, but their modalities are distinct. Specifically, Sequence modal data can be utilized for weather forecasting, such as short-term predictions based on radar data. Multi-modal data can be employed for weather image translation, such as converting multi-modal satellite data to generate radar data. Single-modal data can be applied to various common scenarios, such as radar image super-resolution and post-processing. To unify the diverse weather data modalities into a general representation, we introduce a weather prompt format that assigns different prompt phrases to various modalities. By leveraging in-context learning, our WeatherGFM achieves a promising in-context ability on both various seen tasks and unseen tasks. The significance of our work can be summarized as:
2 Related Work
@@
Weather understanding and beyond. Over the past decade, machine learning techniques (Zhang et al., 2022; 2023; 2024; Wang et al., 2024a) applied in vision tasks have continuously drawn attention within the field of weather and climate. Numerous data-driven machine learning models have been proposed to address classical tasks in weather understanding (Veillette et al., 2020), such as forecasting, super-resolution, image translation, and post-processing. Weather forecasting (Bi et al., 2023a)) aims to predict future observations from past data. Weather super-resolution tasks, i.e., weather downscaling,  Chen et al. (2022) focus on recovering high-resolution data from low-resolution observations. Weather image translation tasks (Stock et al., 2024) involves converting existing observational data into desired target modalities, such as transforming satellite observations into ground-based weather radar data. Post-processing tasks seek to enhance existing model results, such as bias correction and deblurring (Gong et al., 2024a; b).
Despite significant advancements, current methods often rely on specialized datasets and customized single-task models for certain scenarios. Consequently, single-task models struggle to exhibit strong generalization abilities and fail to capture the interconnections between diverse tasks, which hinders the establishment of simulations for the Earth system.
@@
Weather foundation model. The rise of foundation models in Natural Language Processing and computer vision has sparked interest in their application for weather and climate. Large foundation models, enhanced through pre-training, improve the generalization of AI climate models and can be fine-tuned for specific tasks.  Pathak et al. (2022) proposed FourCastNet, a climate pre-trained model using Vision Transformer for high-resolution predictions and rapid inference through self-supervised pre-training and autoregressive fine-tuning. Pangu-Weather (Bi et al., 2023a) utilizes a 3D Earth-specific Transformer for accurate global predictions. FengWu series (Chen et al., 2023a; Han et al., 2024) solve the forecast problem from a multi-modal and multi-task perspective. ClimaX (Nguyen et al., 2023) introduces supervised pre-training to weather prediction, offering flexibility for diverse forecasting tasks. W-MAE (Man et al., 2023) employs unsupervised training via a Masked Auto-Encoder, while MetePFL (Chen et al., 2023b) and FedWing (Chen et al., 2023c)advocate for prompt-based federated learning to enhance collaborative training while protecting privacy. Recent advancements, such as OceanGPT (Bi et al., 2023b), extend LLM capabilities to ocean-related tasks, and ClimateBERT (Webersinke et al., 2021) focuses on processing climate-related texts from extensive sources.
@@
Visual in-context learning. In recent advancements, visual in-context learning has emerged as a promising research area, inspired by the success of language models like GPT-3  (Brown, 2020). These models adapt to various NLP tasks using prompts or in-context examples without extensive retraining. Similarly, in the vision domain, models such as MAE-VQGAN (Hojel et al., 2024) and Painter  (Wang et al., 2023b) have begun exploring in-context learning. However, challenges persist, especially in low-level tasks requiring detailed pixel manipulation. To address this, PromptGIP (Liu et al., 2023b) and GenLV have incorporated in-context learning concepts into their designs to unify low-level vision tasks with diverse input and output modalities, aiming to develop generalist models. Vision-language models (Liu et al., 2024a; b; Zhao et al., 2024a; b) like Unified-IO (Lu et al., 2022) and Unified-IO 2 (Lu et al., 2024) have made significant progress in integrating multiple tasks, highlighting the potential for unified approaches across modalities. Additionally, compositional visual reasoning, exemplified by Visual Programming (Gupta & Kembhavi, 2023), aligns with in-context learning goals by emphasizing visual task synthesis. ViperGPT (Surís et al., 2023) further demonstrates foundational models for visual reasoning, employing computational techniques similar to our objectives, though without relying on programmatic inputs. These collective efforts pave the way for more sophisticated and versatile visual in-context learning frameworks.
@@ 3 Method
As shown in Figure 1, several key weather understanding tasks can be framed using different types of input and output data. For instance, the weather spatial super-resolution (SR) task generates a high-resolution image $x_{HR}$ from a low-resolution image $x_{LR}$, while weather temporal super-resolution predicts a high-resolution image $x^{t}_{HR}$ based on two consecutive observed input images $x^{t-1}_{LR}$ and $x^{t+1}_{LR}$. Weather forecasting relies on a sequence of observed data points $\{x^{1},x^{2},\ldots,x^{t}\}$ to predict future data points $\{x^{t+1},x^{t+2},\ldots\}$. the image translation task focuses on converting an input image from one modality (e.g., satellite image) to another modality (e.g., radar image). Formally, we can represent these tasks as projections from the source input data $X_{S}$ to the target output data $X_{T}$:
@@ 3.2 WeatherGFM: Weather Generalist Foundation Model
Weather in-context learning. Inspired by the success of in-context learning in large language models (Dong et al., 2022) and vision foundation models (Wang et al., 2023a), we propose to unify the weather understanding problem as the visual prompting question-answer paradigm, as illustrated in Eq. 2. Specifically, given a visual question-answer prompt pair $(P_{in},P_{out})$ as a task-guided prompt and a query input $X_{in}$, the model is expected to perceive the context of the prompt (i.e., what task it represents). Consequently, the model can perform the corresponding operations on the query with the prompt. This process can be formulated as follows:
$X_{out}=F_{\tau}(X_{in},P_{in},P_{out};\theta);$ (2)
where $F_{\tau}$ represents a universal foundation model parameterized by $\theta$. We can determine what task will be performed on the query sample by selecting the task-specific prompt.
@@
Mixed-modal mask modeling. Upon redefining the output spaces of the aforementioned representative vision tasks, it is observed that both the input and output of these tasks are in the form of images as transformers-based architectures could provide flexibility by treating the image-like data as a set of tokens. Therefore, we build the WeatherGFM architecture on Vision Transformers (ViT) and propose a mixed-modal masked image modeling (MMIM) pipeline to train multiple weather understanding tasks as shown in Figure 3. Our WeatherGFM comprises three main elements: the format for input data, the architectural design, and loss function.
@@
Loss Function: After generate the output $X_{out}$, optimization objectives are as follows:
$\mathrm{L}_{\theta}(X_{out},X_{T})$ $=\left\|X_{out}-X_{T}\right\|^{2}.$ (5)
where we use MSE (mean square error) loss to train the weather generalist foundation model.
@@ 4 Experiments
To show the versatility of our proposed method, we incorporate up to 10 tasks including diverse weather forecasting, weather super-resolution, weather image translation and weather post-processing tasks into our experiments.
SEVIR. The Storm EVent ImageRy dataset (SEVIR) (Veillette et al., 2020) is a spatiotemporally aligned dataset that contains over 10,000 weather events represented by five spatially and temporally aligned sensors. These sensors consist of three channels (C02, C09, C13) from the GOES-16 satellite, one NEXRAD derived vertically integrated liquid (VIL) mosaics variable, and lighting detections from the GOES GLM sensor. Each SEVIR event spans 4 hours with 5-minute intervals, sampled randomly (with oversampling of events with moderate and high precipitation) using the NOAA Storm Event Database. The SEVIR benchmark supports scientific research on multiple meteorological applications, including future prediction, image-to-image translation, super-resolution, etc. A detailed description of these tasks can be found in Appendix A. In our task, given that the WeatherGFM has the ability to generate images of diverse modalities, we uniformly sample the resolution of images from different modalities to 256×256. Moreover, we perform filtering on the events within the SEVIR dataset and pick out those events that include both the three channels of the GOES-16 satellite and the one variable derived from weather radar. Ultimately, the dataset we utilize comprises 11,508 events with four distinct sensing modalities. Among them, 11,308 events are selected as the training set, while 100 events are designated as the validation set and 100 events are designated as the test set. Each SEVIR event covers a period of 4 hours with increments of 5 minutes, which implies that there are 49 images for each modality within a single event. Consequently, the training set contains a total of 2.2M images, while the validate/test set has a total of 19.6K images. Regarding the demarcation of the training set and the test set for each task, we provide a detailed introduction in Appendix A.
@@ 4.2 Implementation and Evaluation
To count the Hits (truth=1, pred=1), Misses (truth=1, pred=0) and F.Alarms (truth=0, pred=1), the prediction and the ground-truth are normalized using mean-variance normalization and binarized at different thresholds. For radar output tasks, we have established thresholds at [16, 74, 133, 160, 181, 219]. GEOS-visible output tasks are assigned thresholds of [2000, 3200, 4400, 5600, 6800]. The GEOS-IR107 output tasks operate with thresholds set to [-6000, -4000, 0, 2000]. Lastly, the GEOS-IR069 output task employs thresholds of [-4000, -5000, -6000, -7000].

### v2

@@ Abstract
The Earth’s weather system involves intricate weather data modalities and diverse weather understanding tasks, which hold significant value to human life. Existing data-driven models focus on single weather understanding tasks (e.g., weather forecasting). While these models have achieved promising results, they fail to tackle various complex tasks within a single and unified model. Moreover, the paradigm that relies on limited real observations for a single scenario hinders the model’s performance upper bound. Inspired by the in-context learning paradigm from visual foundation models and large language models, in this paper, we introduce the first generalist weather generalist foundation model (WeatherGFM) to address weather understanding tasks in a unified manner. Specifically, we first unify the representation and definition for diverse weather understanding tasks. Subsequently, we design weather prompt formats to handle different weather data modalities, including single, multiple, and temporal modalities. Finally, we adopt a visual prompting question-answering paradigm for the training of unified weather understanding tasks. Extensive experiments indicate that our WeatherGFM can effectively handle up to ten weather understanding tasks, including weather forecasting, super-resolution, weather image translation, and post-processing. Our method also showcases generalization ability on unseen tasks.
@@ 1 Introduction
Modeling Earth weather systems involves a series of complex subprocesses that are intended to transform intricate Earth observation data into applications like weather forecasting (Chen et al., 2023a; Bi et al., 2023), downscaling (Chen et al., 2022), assimilation (Huang et al., 2024), retrieval (Liu et al., 2011), and bias correction (Gong et al., 2024). During the past decade, many data-driven machine learning methods have been investigated for various weather understanding tasks and delivering desirable performance on specific tasks. For example, recent studies using large-scale training data (e.g., ERA5 reanalysis data (Hersbach et al., 2020)) have exceeded the accuracy of conventional numerical weather forecasts. However, current weather foundational models face challenges regarding generalizability and data scale limitations. On the one hand, the Earth observation system consists of a variety of observation devices, such as satellites, radar, and weather stations, which produce diverse modalities of data. Consequently, designing a specific model for a single-task scenario is highly complex, time-consuming, and labor-intensive. On the other hand, large-scale data in fields such as computer vision can be obtained at a low cost, whereas weather understanding tasks face an intrinsic bottleneck in data scale due to restrictions on individual scenes and single observation devices as shown in Table 1. For instance, local short-term precipitation forecasting models can only utilize a finite range of observational data.
A significant trend in AI research is the development of foundation models, shifting towards large-scale pre-training and in-context learning. This paradigm enables unified processing of a multitude of complex tasks and generalization to unseen tasks. For example, large language models (LLMs) can perform a variety of language-centric tasks (e.g., sentiment analysis, question answering and machine translation) by combining language input-output examples with new query inputs (prompts) without optimizing model parameters (Brown, 2020). Similarly, vision foundation models (Wang et al., 2023b; Liu et al., 2023b; Chen et al., 2024b) employ visual prompts with query inputs to carry out diverse image-centric tasks, such as semantic segmentation, depth estimation, and image restoration. These studies highlight the significant potential of generalist foundational models.
@@
In this paper, we first propose a weather generalist foundation model, WeatherGFM, to uniformly address a variety of complex weather understanding tasks and data modalities. Unlike prior studies that focused on weather forecasting, our proposed method can expand the task scope to weather forecasting, weather super-resolution (i.e., weather downscaling) (Veillette et al., 2020), weather image translation (similar to retrieval in weather) (Veillette et al., 2020), and post-processing (Gong et al., 2024). These tasks all belong to the domain of weather understanding, but their modalities are distinct. Specifically, Sequence modal data can be utilized for weather forecasting, such as short-term predictions based on radar data. Multi-modal data can be employed for weather image translation, such as converting multi-modal satellite data to generate radar data. Single-modal data can be applied to various common scenarios, such as radar image super-resolution and post-processing. To unify the diverse weather data modalities into a general representation, we introduce a weather prompt format that assigns different prompt phrases to various modalities. By leveraging in-context learning, our WeatherGFM achieves a promising in-context ability on both various seen tasks and unseen tasks. The significance of our work can be summarized as:
2 Related Work
@@
Computer Vision HAT (Chen et al., 2023b) Low-cost Image super-resolution (SR)
@@
Cascast (Gong et al., 2024) High-cost Post-processing
Climax (Nguyen et al., 2023) High-cost Weather forecasting, Super-resolution Requires fine-tuning
@@
Weather understanding and beyond. Over the past decade, machine learning techniques have consistently attracted attention in the field of weather and climate. Numerous data-driven machine learning models have been proposed to address classical tasks in weather understanding (Veillette et al., 2020), such as forecasting, super-resolution, image translation, and post-processing. Weather forecasting (Bi et al., 2023)) aims to predict future observations from past data. Weather super-resolution tasks, i.e., weather downscaling,  Chen et al. (2022) focus on recovering high-resolution data from low-resolution observations. Weather image translation tasks (Stock et al., 2024) involves converting existing observational data into desired target modalities, such as transforming satellite observations into ground-based weather radar data. Post-processing tasks seek to enhance existing model results, such as bias correction and deblurring (Gong et al., 2024). Despite significant advancements, current methods often rely on specialized datasets and customized single-task models for certain scenarios. Consequently, single-task models struggle to exhibit strong generalization abilities and fail to capture the interconnections between diverse tasks, which hinders the establishment of simulations for the Earth system.
@@
Weather foundation model. The rise of foundation models (Liu et al., 2024; Zhao et al., 2024a; b) in Natural Language Processing and computer vision has sparked interest in their application for weather and climate. Large foundation models, enhanced through pre-training, improve the generalization of AI climate models and can be fine-tuned for specific tasks.  Pathak et al. (2022) proposed FourCastNet, a climate pre-trained model using Vision Transformer for high-resolution predictions and rapid inference through self-supervised pre-training and autoregressive fine-tuning. Pangu-Weather (Bi et al., 2023) utilizes a 3D Earth-specific Transformer for accurate global predictions. ClimaX (Nguyen et al., 2023) introduces supervised pre-training to weather prediction, offering flexibility for diverse forecasting tasks. A pre-training foundation model usually requires mask modeling for pre-training and then undergoes fine-tuning on specific tasks, such as fine-tuning the pre-trained model on weather forecasting, remote sensing classification and segmentation tasks Bodnar et al. (2024); Cong et al. (2022); Noman et al. (2024); Li et al. (2024).
@@
Visual in-context learning. In recent advancements, visual in-context learning has emerged as a promising research area, inspired by the success of language models like GPT-3  (Brown, 2020). These models adapt to various NLP tasks using prompts or in-context examples without extensive retraining. Similarly, in the vision domain, models such as MAE-VQGAN (Hojel et al., 2024) and Painter  (Wang et al., 2023b) have begun exploring in-context learning. However, challenges persist, especially in low-level tasks requiring detailed pixel manipulation. To address this, PromptGI (Liu et al., 2023b) and GenLV have incorporated in-context learning concepts into their designs to unify low-level vision tasks with diverse input and output modalities, aiming to develop generalist models. Vision-language models like Unified-IO (Lu et al., 2022) and Unified-IO 2 (Lu et al., 2024) have made significant progress in integrating multiple tasks, highlighting the potential for unified approaches across modalities. Additionally, compositional visual reasoning, exemplified by Visual Programming (Gupta & Kembhavi, 2023), aligns with in-context learning goals by emphasizing visual task synthesis. ViperGPT (Surís et al., 2023) further demonstrates foundational models for visual reasoning, employing computational techniques similar to our objectives, though without relying on programmatic inputs. These collective efforts pave the way for more sophisticated and versatile visual in-context learning frameworks.
@@ 3 Method
As shown in Figure 1, several key weather understanding tasks can be framed using different types of input and output data. For instance, the weather spatial super-resolution (SR) task generates a high-resolution image $x_{HR}$ from a low-resolution image $x_{LR}$, while weather temporal super-resolution predicts a high-resolution image $x^{t}_{HR}$ based on two consecutive observed input images $x^{t-1}_{LR}$ and $x^{t+1}_{LR}$, where $t$ represents a particular moment in time. The weather temporal super-resolution task aims to restore the missing observed data in time $t$. Weather forecasting relies on a sequence of observed data points $\{x^{1},x^{2},\ldots,x^{t}\}$ that are gathered over the past $t$ time steps. These observed data points serve as condition, enabling the prediction of future data points such as points $\{x^{t+1},x^{t+2},\ldots\}$. The image translation task focuses on converting an input image from one modality (e.g., satellite image) to another modality (e.g., radar image). Formally, we can represent these tasks as projections from the source input data $X_{S}$ to the target output data $X_{T}$:
@@ 3.2 WeatherGFM: Weather Generalist Foundation Model
Weather in-context learning. Inspired by the success of in-context learning in large language models (Dong et al., 2022) and vision foundation models (Wang et al., 2023a), we propose to unify the weather understanding problem as the visual prompting question-answer paradigm, as illustrated in Eq. 3. Specifically, given a visual question-answer prompt pair $(P_{in},P_{target})$ as a task-guided prompt and a query input $X_{in}$, the model is expected to perceive the context of the prompt (i.e., what task it represents). Consequently, the model can perform the corresponding operations on the query with the prompt. This process can be formulated as follows:
$X_{target}=F_{\tau}(P_{in},P_{target},X_{in};\theta);$ (2)
where $F_{\tau}$ represents a universal foundation model parameterized by $\theta$. $P_{in}$ and $P_{target}$ denotes the input and target of task prompts. We can determine what task will be performed on the input $X_{in}$ by selecting the task-specific prompt $P_{in}$ and $P_{target}$, and then obtain the target $X_{target}$ for the corresponding task through the model $F_{\tau}$.
@@
Mixed-modal mask modeling. Upon redefining the output spaces of the aforementioned representative vision tasks, it is observed that both the input and output of these tasks are in the form of images as transformers-based architectures could provide flexibility by treating the image-like data as a set of tokens. Therefore, we build the WeatherGFM architecture on Vision Transformers (ViT) and propose a mixed-modal masked image modeling (MMIM) pipeline to train multiple weather understanding tasks as shown in Figure 3. Inspired by the concept of Visual Question Answering Wang et al. (2023a); Liu et al. (2023b); Chen et al. (2024a), we introduce mixed-modality masking on various weather modalities for visual question-and-answer modeling in weather understanding tasks. This process can be formulated as follows:
$P^{{}^{\prime}}_{target},X^{{}^{\prime}}_{target}=F_{\tau}(P_{in},M(P_{target}),X_{in},M(X_{target});\theta);$ (3)
where we randomly conduct mask operation $M$ on the prompt target $P_{target}$ as well as the ground truth $X_{target}$ according to the mask ratio. Meanwhile, the prompt input $P_{in}$ and the input query $X_{in}$ will be retained entirely. $P^{{}^{\prime}}_{target}$ and $X^{{}^{\prime}}_{target}$ represent the predicted target output of model $F_{\tau}$. The optimization objectives are as follows:
$\mathrm{L}^{total}_{\theta}$ $=\mathrm{L}_{2}(P^{{}^{\prime}}_{target},P_{target})+\mathrm{L}_{2}(X^{{}^{\prime}}_{target},X_{target}).$ (4)
where we use MSE (mean square error) loss $\mathrm{L_{2}}$ to train the weather generalist foundation model. In the inference stage, we keep the $P_{in}$, $P_{target}$, and $X_{in}$ intact while the target image is fully masked. This target full masking strategy allows generalist foundation models to generate the corresponding target through a visual question-and-answer format. Our WeatherGFM comprises two main elements: the format for input data and the architectural design.
@@ 4 Experiments
We incorporate up to 10 tasks including diverse weather forecasting, weather super-resolution, weather image translation and weather post-processing tasks into our experiments.
SEVIR. The Storm EVent ImageRy dataset (SEVIR) (Veillette et al., 2020) is a spatiotemporally aligned dataset that contains over 10,000 weather events represented by five spatially and temporally aligned sensors. These sensors consist of three channels (C02, C09, C13) from the GOES-16 satellite, one NEXRAD derived vertically integrated liquid (VIL) mosaic variable, and lighting detections from the GOES GLM sensor. Each SEVIR event spans 4 hours with 5-minute intervals, sampled randomly (with oversampling of events with moderate and high precipitation) using the NOAA Storm Event Database. In our task, we uniformly resize the resolution of images from different modalities to 256×256. Moreover, we filter the events within the SEVIR dataset and pick out those events that include both the three channels of the GOES-16 satellite and the one variable derived from weather radar. Ultimately, the dataset we utilize comprises 11,508 events with four distinct sensing modalities. Among them, 11,308 events are selected as the training set, while 100 events are designated as the validation set and 100 events are designated as the test set. Consequently, the training set contains a total of 2.2M images, while the validate/test set has a total of 19.6K images. We provide a detailed introduction in Appendix A.
@@ 4.2 Implementation and Evaluation
To count the Hits (truth=1, pred=1), Misses (truth=1, pred=0) and F.Alarms (truth=0, pred=1), the prediction and the ground-truth are normalized using mean-variance normalization and binarized at different thresholds. Following SEVIR (Veillette et al., 2020), for radar output tasks, we have established thresholds at [16, 74, 133, 160, 181, 219]. GEOS-visible output tasks are assigned thresholds of [2000, 3200, 4400, 5600, 6800]. The GEOS-IR107 output tasks operate with thresholds set to [-6000, -4000, 0, 2000]. Lastly, the GEOS-IR069 output task employs thresholds of [-4000, -5000, -6000, -7000].